OpenStudy (albert0898):

*** WILL GIVE MEDAL! *** Which value of x is not in the domain of the function defined by y = tan x? (1) 2pi/3 (2) pi/2 (3) pi (4) pi/3

OpenStudy (solomonzelman):

tan(x) = sin(x) / cos(x) you know that, correct?

OpenStudy (albert0898):

Yes

OpenStudy (solomonzelman):

whatever value(s) of x that make(s) the denominator ( the cos(x) ) be equal to 0, is/are the value(s) where the function y is undefined.

OpenStudy (solomonzelman):

(( In other words, those values of x that make the denominator be equal to 0 are not in the domain of the function ))

OpenStudy (albert0898):

The answer is pi/2 , but how do I get to that? Still confused

OpenStudy (albert0898):

@SolomonZelman @phi

OpenStudy (solomonzelman):

do you know what \(\normalsize\color{black}{ \displaystyle \cos(\pi/2) }\) is?

OpenStudy (albert0898):

0

OpenStudy (albert0898):

Ohhhhhhhh

OpenStudy (solomonzelman):

any questions ?

OpenStudy (albert0898):

Not at this time, thank you!